OpenStudy (igster1112):

What type of measurement is indicated by each of the following units? 1. g/ml 2. s 3. km 4. g 5. cm³ 6. mm 7. mg 8. L 9. g/cm³

Nnesha (nnesha):

do you know what `ml` standz for ?

OpenStudy (igster1112):

Milliliters

OpenStudy (igster1112):

Ok thank you

Nnesha (nnesha):

\[\huge\rm D=\frac{mass ~\color{ReD}{g}}{Volume ~\color{red}{mL}}\] that's the formula for density where mass should be in grams and volume in mL

OpenStudy (igster1112):

So anything that has a g in in would be mass? Like mg? And g/cm³?

Nnesha (nnesha):

mass can be written in kg or g or milligram g/cm^3 is not just mass it is g `over` cm^3 cm^2 = centimeter cubed (unit for volume ) \[\rm \frac{g}{cm^2} =\frac {mass}{volume }\] is a unit of what ??? mass/volume

OpenStudy (igster1112):

Oh ok I get it